Determine the maximum area of a triangle, in square centimeters, if the sum of its base and its height is 10 cm.
Answer by checkley77(12844)

area=bh/2
area=5*5/2=25/2=12.5 cm^2.
Proof:
5.1*4.9/2=24.99/2=12.495 cm^2.